A marble block of mass 2 kg lying on ice when given a velocity of 6 m/s is stopped by friction in 10 s. then the coefficient of friction is ?

Grade:12

1 Answers

Asaad
6 years ago
u = 6 m/s v = 0t = 10 sμ = ? v = u + at 0 = 6 + 10aa = - 3/5 m/s2Frictional force=ma=μmgma=μmga=μg